OpenStudy (anonymous):

Determine if (3/8, -1) is a solution to 8x-7y=10.

OpenStudy (anonymous):

just substitute x=3/8 and y=-1 into the equation. if it equal to 10, then (3/8,-1) a solution.

OpenStudy (anonymous):

can you show how it is done because math is a bad subject for me and I am very confused

OpenStudy (anonymous):

u hav (3/8,-1) and equation: 8x-7y=10 substitute x and y, u wil hav: 8(3/8)-7(-1)=10. this implies 3+7=10 => 10=10 since the equations holds for (3/8,-1), hence it is a solution for the equation.

OpenStudy (anonymous):

ok so it is a solution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can u please tell me why it won't let me post another question to be helped on